OpenStudy (anonymous):

The length of a running track is 300/400 kilometers. A runner covers two thirds of it. How many kilometers has he covered?

OpenStudy (anonymous):

\[\frac{ 300 }{400 }\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The length of a running track is 300/400 kilometers. A runner covers two thirds of it. How many kilometers has he covered?

OpenStudy (anonymous):

ok so you get your length \[\frac{ 300 }{ 400 } \times \frac{ 2 }{ 3 }\]